# See pages - of Main.pdf

### v1

@@ 1 Piezoelectricity extraction
Figure 1: Microscope image of a representative IDT fabricated on a SrTiO3 thin film.
@@ 1.1 Electrical characterization and impedance extraction
Figure 2: a) Mason equivalent circuit model for the IDT electromechanical response. The typical measured resistance and wideband $\mathrm{Im}[Z]$ are shown in (b) and (c), respectively. (d) Shows the numerically simulated map from capacitance to SrTiO3 permittivity.
@@ 1.2 Mason model and electromechanical coupling
The electromechanical response of the IDTs is better understood using Mason’s equivalent circuit model, in which the total electrical admittance $Y(f)$ is expressed as the sum of a static capacitive contribution and a frequency-dependent motional response[12]:
$Y(f)=i\omega C_{T}+G_{m}(f)+iB_{m}(f),$ (2)
@@
$G_{m}(f)=8K^{2}C_{T}Nf_{0}\left(\frac{\sin X}{X}\right)^{2}\equiv G_{0}\left(\frac{\sin X}{X}\right)^{2},\qquad X=N\pi\frac{|f-f_{0}|}{f_{0}},$ (3)
@@
Both $G_{m}(f)$ and $B_{m}(f)$ are narrow-band features centered around $f_{0}$ with a fractional bandwidth of $\Delta f/f_{0}\approx 0.885/N$. Figure 2(a) shows a circuit representation of the Mason model with addition of parasitic resistance ($R_{s}$) and inductance ($L_{s}$). Fig. 2(b,c) show a typical measured resistance $\text{Re}[Z]$, and wide-band Im[Z][Z], respectively.
@@
As evident from Fig. 2 (b), the parasitic resistance contributes a constant background to the measured response. In the following, we focus on extracting the motional resistance that appears on top of this background. At resonance, the motional resistance is given by
$R_{m}(f_{0})=\Re\{1/Y(f_{0})\}=\frac{G_{0}}{G_{0}^{2}+\left(\omega_{0}C_{T}\right)^{2}}.$ (5)
@@
The motional resistance is governed by the electromechanical coupling coefficient $K^{2}$, which itself is determined by the piezoelectric tensor ($\mathbf{e}$), permittivity ($\mathbf{\epsilon}$) and elasticity ($\mathbf{c}$) tensors of the SrTiO3 film. In bulk acoustic wave transducers, $K^{2}$ is given by $K^{2}=\frac{e_{33}^{2}}{c_{33}\varepsilon_{33}}$  [11]. However, for surface acoustic wave (SAW) transducers, $K^{2}$ does not have a closed-form expression and can be regarded as a function $K^{2}=f(\mathbf{e,\epsilon,c})$. Therefore, we numerically simulate the IDT electromechanical response using 2D-FEM in COMSOL Multiphysics to model the Rayleigh wave and extract the SrTiO3 material properties. The entire number of fingers was simulated to calculate the motional resistance. Since the simulation is two dimensional, the shear waves (out-of-plane in the 2D simulation) is not captured. The required permittivity extraction is extracted using the COMSOL map in Fig 2(d) as discussed earlier in the methods section. The elasticity tensor is taken from literature [2, 13], namely, $c_{11}=337.23$ [GPa], $c_{12}=132.2$ [GPa], $c_{44}=115.34$ [GPa]. In the simulation, the elasticity of SrTiO3 was modeled as a ”cubic” tensor, where the diagonal elements satisfy $c_{11}=c_{22}=c_{33}$, the off-diagonal elements $c_{ij}$ (for $i\neq j$ and $i,j=1,2,3$) are equal, and the shear components satisfy $c_{44}=c_{55}=c_{66}$. This approach shows a good agreement with our acoustic dispersion results presented in the main text. The only free parameter is the piezoelectric tensor $\mathbf{e}$, which is adjusted to fit the amplitude of the measured motional resistance.
@@
For Rayleigh-wave excitation, the dominant piezoelectric tensor components are $e_{33}$ and $e_{31}$, consistent with the predominantly in-plane displacement of Ti ions in the SrTiO3 thin films as discussed in section 1.4. The $e_{31}$ component is negative and can be physically understood as the transverse contraction of the unit cell accompanying expansion along the polar axis. For ABO3 perovskites, $|e_{31}|$ is typically at least one order of magnitude smaller than $e_{33}$ [21]. Density functional theory calculations for SrTiO3 suggest even smaller values of $e_{31}$ [6]. In our analysis, we therefore assume $e_{31}=-0.1\,e_{33}$, as an upper bound estimation for $e_{31}$ (lower bound for $e_{33}$) based on related perovskite materials. In summary, the extraction methodology for $e_{33,\mathrm{eff}}$ is as follows:
@@ 1.3 $e_{15}$ extraction based on BAW
The transducer can also be considered as a lateral field excitation (LFE) transducer with parallel electrodes configuration. In this case, the analysis is simpler, the electromechanical coupling coefficient for thickness shear wave excitation is given by [11, 5]
$K_{15,\mathrm{eff}}=\frac{e_{15,\mathrm{eff}}}{\sqrt{c_{55}^{E}\varepsilon_{11}^{T}}}.$ (6)
@@
$K_{15,\mathrm{eff}}^{2}=\left(\frac{\pi}{2}\right)^{2}\frac{f_{p}-f_{s}}{f_{p}},$ (7)
where $f_{p}$ is the parallel resonance frequency and $f_{s}$ is the series resonance frequency shown in Fig. 3.
@@ 1.4 Acoustic wave excitation dependency on piezoelectric tensor
To understand the dependency of Rayleigh wave excitation on different piezoelectric tensor components, we solve the governing equations for acoustic wave as an eigenvalue problem to obtain the acoustic modes’ profiles and their corresponding eigenfrequencies (velocities) in COMSOL Multiphysics, namely [23],
$\nabla_{s}\left(\mathbf{c}\left[\nabla_{s}\right]^{T}\mathbf{u}+\mathbf{e}^{T}[\nabla]^{T}\phi\right)+\rho\omega^{2}\mathbf{u}$ $=0,$ (8)
@@
where $\mathbf{u}$ is the displacement vector, $\phi$ is the electric potential, $\mathbf{c}$ is the elasticity tensor, $\mathbf{e}$ is the piezoelectric stress-tensor, $\boldsymbol{\varepsilon}$ is the permittivity tensor and $\rho$ is the density. For a given acoustic mode, the magnitude of the gap between dispersion curves calculated with and without piezoelectric coupling provides a direct measure of how efficiently that mode can be electrically excited. Piezoelectric coupling modifies the acoustic phase velocity; by reciprocity, this implies that an electric field applied along the corresponding crystallographic direction can efficiently drive the same acoustic mode. Guided by this principle, we perform single-domain simulations both including piezoelectric coupling, implemented via free-floating electric potential boundary conditions (BCs), and excluding piezoelectric coupling, enforced through grounded potential BCs. This procedure enables the systematic construction of slowness curves (inverse phase velocity) as a function of propagation direction and allows the contribution of individual piezoelectric tensor components to be isolated. Figure 4a shows the slowness curves of the Rayleigh wave under free and grounded potential BCs, considering only the $e_{33}$ component. It follows that the electromechanical coupling can be calculated from the slowness curves as [23]
$K^{2}=2\times\left(1-\frac{v_{\mathrm{grounded}}}{v_{\mathrm{free}}}\right)\times 100\%.$ (9)
@@
Figure 4: (a) Slowness curves for Rayleigh wave with and without piezoelectric coupling accounting only $e_{33}$ component. (b) Rayleigh wave electromechanical coupling for different piezoelectric tensor elements.
Figure 4(b) shows the dependency of the Rayleigh wave excitation on the piezoelectric coefficient $e_{33}$. The wave can be excited efficiently in the c-oriented domain but not in the a-oriented domain. It can be seen that the $e_{31}$ component slightly enhances the excitation, while the $e_{15}$ component doesn’t contribute to the Rayleigh wave excitation with 0 and 90 degrees IDTs. Similar interpretations can be drawn by merely looking at the rotated piezoelectric tensor. As a reminder, the piezoelectric tensor for tetragonal c-axis (Ti atom is out of the wafer’s plane) SrTiO3 is given by
$e]=\begin{bmatrix}0&0&0&0&e_{15}&0\\ 0&0&0&e_{15}&0&0\\ e_{31}&e_{31}&e_{33}&0&0&0\end{bmatrix}$. (10)
@@
$e]_{a}^{0}=\begin{array}[]{cccccc}S_{1}\quad S_{2}\quad S_{3}\quad S_{4}\quad S_{5}\quad S_{6}\\ \left[\begin{array}[]{cccccc}0&0&0&0&0&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{15}\\ e_{31}&e_{33}&e_{31}&0&0&0\\ 0&0&0&e_{15}&0&0\\ \end{array}\right]\\ \end{array}\begin{gathered}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}E_{x}\\ E_{y}\\ E_{z}\end{gathered}$ (11)
@@
$e]_{c}^{90}=\begin{array}[]{cccccc}S_{1}\quad S_{2}\quad S_{3}\quad S_{4}\quad S_{5}\quad S_{6}\\ \left[\begin{array}[]{cccccc}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{33}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{31}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{31}&0&0&0\\ 0&0&0&0&0&e_{15}\\ 0&0&0&0&e_{15}&0\\ \end{array}\right]\\ \end{array}\begin{gathered}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}E_{x}\\ E_{y}\\ E_{z}\end{gathered}$ (12)
For a dominantly x-oriented electric field across $e]_{a}^{0}$ the $e_{15}$ now is at 1,6 position which couples x-field to shear x-y stress, consistent with the mode profile shown in Fig. 5a. Similarly, for x-oriented electric field across $e]_{c}^{90}$, the $e_{33}$ now is at 1,1 position which couples x-field to x -longitudinal stress, and $e_{31}$ at the 1,3 position which couples x-field to z- transverse stresses. This matches the elliptical motion that builds the Rayleigh wave.
Figure 5: Simulated mode profiles with its corresponding driving domain for (a) Rayleigh wave and (b) Bulk shear wave.
@@ 1.5 Single domain piezoelectric coefficient of twin-domain structure
Starting from the electromechanical coupling coefficient definition [3]
$K=\frac{U_{m}}{\sqrt{U_{e}U_{d}}},$ (13)
@@
$U_{m}=\frac{1}{2}\left[\int_{\Omega_{a}}\left(\frac{e_{33}^{(a)}}{c_{33}^{E(a)}}\right)E_{3}T_{3}dV+\int_{\Omega_{c}}\left(\frac{e_{33}^{(c)}}{c_{33}^{E(c)}}\right)E_{3}T_{3}dV\right],$ (14)
$U_{e}$ $=\frac{1}{2}\int_{\Omega}\frac{1}{c_{33}^{E}}T_{3}^{2}dV,$ (15)
$U_{d}$ $=\frac{1}{2}\int_{\Omega}\varepsilon_{33}^{T}E_{3}^{2}dV.$ (16)
@@
$\operatorname{Vol}\left(\Omega_{a}\right)=\operatorname{Vol}\left(\Omega_{c}\right)=\frac{1}{2}\operatorname{Vol}(\Omega).$ (17)
@@
$U_{m}=\frac{1}{2}\left[\int_{\Omega_{x}}\left(\frac{e_{15}^{(x)}}{c_{55}^{E(x)}}\right)\cos(45^{\circ})^{2}E_{1}T_{5}dV+\int_{\Omega_{y}}\left(\frac{e_{15}^{(y)}}{c_{55}^{E(y)}}\right)\cos(45^{\circ})^{2}E_{1}T_{5}dV\right].$ (19)
@@
$U_{e}$ $=\frac{1}{2}\int_{\Omega}\frac{1}{c_{55}^{E}}T_{5}^{2}dV,$ (20)
$U_{d}$ $=\frac{1}{2}\int_{\Omega}\varepsilon_{55}^{T}E_{1}^{2}dV.$ (21)

### v2

# Supplementary Materials for: Cryogenic piezoelectric effects in thin film strontium titanate devices

@@
See pages - of Main.pdf
Supplementary Materials for
Cryogenic piezoelectric effects in thin film
strontium titanate devices
Ahmed Khalil et al.
Corresponding author. Email: Christian.Haffner@imec.be
This PDF file includes:
Supplementary Text
Figures S1 to S9
Tables S1 to S2
@@ 1 Piezoelectricity extraction
Fig. S1: IDT on SrTiO3 thin film. Microscope image of a representative IDT fabricated on a SrTiO3 thin film.
@@ 1.1 Electrical characterization and impedance extraction
Fig. S2: Mason model and permittivity extraction. (a) Mason equivalent circuit model for the IDT electromechanical response. The typical measured resistance and wideband $\mathrm{Im}[Z]$ are shown in (b) and (c), respectively. (d) Shows the numerically simulated map from capacitance to SrTiO3 permittivity.
@@ 1.2 Mason model and electromechanical coupling
The electromechanical response of the IDTs is better understood using Mason’s equivalent circuit model, in which the total electrical admittance $Y(f)$ is expressed as the sum of a static capacitive contribution and a frequency-dependent motional response  (?):
$Y(f)=i\omega C_{T}+G_{m}(f)+iB_{m}(f),$ (S2)
@@
$G_{m}(f)=8K^{2}C_{T}Nf_{0}\left(\frac{\sin X}{X}\right)^{2}\equiv G_{0}\left(\frac{\sin X}{X}\right)^{2},\qquad X=N\pi\frac{|f-f_{0}|}{f_{0}},$ (S3)
@@
Both $G_{m}(f)$ and $B_{m}(f)$ are narrow-band features centered around $f_{0}$ with a fractional bandwidth of $\Delta f/f_{0}\approx 0.885/N$. Fig. S2(a) shows a circuit representation of the Mason model with addition of parasitic resistance ($R_{s}$) and inductance ($L_{s}$). Fig. S2(b,c) show a typical measured resistance $\text{Re}[Z]$, and wide-band Im[Z][Z], respectively.
@@
As evident from Fig. S2 (b), the parasitic resistance contributes a constant background to the measured response. In the following, we focus on extracting the motional resistance that appears on top of this background. At resonance, the motional resistance is given by
$R_{m}(f_{0})=\Re\{1/Y(f_{0})\}=\frac{G_{0}}{G_{0}^{2}+\left(\omega_{0}C_{T}\right)^{2}}.$ (S5)
@@
The motional resistance is governed by the electromechanical coupling coefficient $K^{2}$, which itself is determined by the piezoelectric tensor ($\mathbf{e}$), permittivity ($\mathbf{\epsilon}$) and elasticity ($\mathbf{c}$) tensors of the SrTiO3 film. In bulk acoustic wave transducers, $K^{2}$ is given by $K^{2}=\frac{e_{33}^{2}}{c_{33}\varepsilon_{33}}$   (?). However, for surface acoustic wave (SAW) transducers, $K^{2}$ does not have a closed-form expression and can be regarded as a function $K^{2}=f(\mathbf{e,\epsilon,c})$. Therefore, we numerically simulate the IDT electromechanical response using 2D-FEM in COMSOL Multiphysics to model the Rayleigh wave and extract the SrTiO3 material properties. The entire number of fingers was simulated to calculate the motional resistance. Since the simulation is two dimensional, the shear waves (out-of-plane in the 2D simulation) is not captured. The required permittivity extraction is extracted using the COMSOL map in Fig. S2(d) as discussed earlier in the methods section. The elasticity tensor is taken from literature  (?, ?), namely, $c_{11}=337.23$ [GPa], $c_{12}=132.2$ [GPa], $c_{44}=115.34$ [GPa]. In the simulation, the elasticity of SrTiO3 was modeled as a ”cubic” tensor, where the diagonal elements satisfy $c_{11}=c_{22}=c_{33}$, the off-diagonal elements $c_{ij}$ (for $i\neq j$ and $i,j=1,2,3$) are equal, and the shear components satisfy $c_{44}=c_{55}=c_{66}$. This approach shows a good agreement with our acoustic dispersion results presented in the main text. The only free parameter is the piezoelectric tensor $\mathbf{e}$, which is adjusted to fit the amplitude of the measured motional resistance.
@@
For Rayleigh-wave excitation, the dominant piezoelectric tensor components are $e_{33}$ and $e_{31}$, consistent with the predominantly in-plane displacement of Ti ions in the SrTiO3 thin films as discussed in section 1.4. The $e_{31}$ component is negative and can be physically understood as the transverse contraction of the unit cell accompanying expansion along the polar axis. For ABO3 perovskites, $|e_{31}|$ is typically at least one order of magnitude smaller than $e_{33}$  (?). Density functional theory calculations for SrTiO3 suggest even smaller values of $e_{31}$  (?). In our analysis, we therefore assume $e_{31}=-0.1\,e_{33}$, as an upper bound estimation for $e_{31}$ (lower bound for $e_{33}$) based on related perovskite materials. In summary, the extraction methodology for $e_{33,\mathrm{eff}}$ is as follows:
@@ 1.3 $e_{15}$ extraction based on BAW
The transducer can also be considered as a lateral field excitation (LFE) transducer with parallel electrodes configuration. In this case, the analysis is simpler, the electromechanical coupling coefficient for thickness shear wave excitation is given by (?, ?)
$K_{15,\mathrm{eff}}=\frac{e_{15,\mathrm{eff}}}{\sqrt{c_{55}^{E}\varepsilon_{11}^{T}}}.$ (S6)
@@
$K_{15,\mathrm{eff}}^{2}=\left(\frac{\pi}{2}\right)^{2}\frac{f_{p}-f_{s}}{f_{p}},$ (S7)
where $f_{p}$ is the parallel resonance frequency and $f_{s}$ is the series resonance frequency shown in Fig. S3.
@@ 1.4 Acoustic wave excitation dependency on piezoelectric tensor
To understand the dependency of Rayleigh wave excitation on different piezoelectric tensor components, we solve the governing equations for acoustic wave as an eigenvalue problem to obtain the acoustic modes’ profiles and their corresponding eigenfrequencies (velocities) in COMSOL Multiphysics, namely (?),
$\nabla_{s}\left(\mathbf{c}\left[\nabla_{s}\right]^{T}\mathbf{u}+\mathbf{e}^{T}[\nabla]^{T}\phi\right)+\rho\omega^{2}\mathbf{u}$ $=0,$ (S8)
@@
where $\mathbf{u}$ is the displacement vector, $\phi$ is the electric potential, $\mathbf{c}$ is the elasticity tensor, $\mathbf{e}$ is the piezoelectric stress-tensor, $\boldsymbol{\varepsilon}$ is the permittivity tensor and $\rho$ is the density. For a given acoustic mode, the magnitude of the gap between dispersion curves calculated with and without piezoelectric coupling provides a direct measure of how efficiently that mode can be electrically excited. Piezoelectric coupling modifies the acoustic phase velocity; by reciprocity, this implies that an electric field applied along the corresponding crystallographic direction can efficiently drive the same acoustic mode. Guided by this principle, we perform single-domain simulations both including piezoelectric coupling, implemented via free-floating electric potential boundary conditions (BCs), and excluding piezoelectric coupling, enforced through grounded potential BCs. This procedure enables the systematic construction of slowness curves (inverse phase velocity) as a function of propagation direction and allows the contribution of individual piezoelectric tensor components to be isolated. Fig. S4a shows the slowness curves of the Rayleigh wave under free and grounded potential BCs, considering only the $e_{33}$ component. It follows that the electromechanical coupling can be calculated from the slowness curves as (?)
$K^{2}=2\times\left(1-\frac{v_{\mathrm{grounded}}}{v_{\mathrm{free}}}\right)\times 100\%.$ (S9)
@@
Fig. S4: Rayleigh wave slowness curves and coupling. (a) Slowness curves for Rayleigh wave with and without piezoelectric coupling accounting only $e_{33}$ component. (b) Rayleigh wave electromechanical coupling for different piezoelectric tensor elements.
Fig. S4(b) shows the dependency of the Rayleigh wave excitation on the piezoelectric coefficient $e_{33}$. The wave can be excited efficiently in the c-oriented domain but not in the a-oriented domain. It can be seen that the $e_{31}$ component slightly enhances the excitation, while the $e_{15}$ component doesn’t contribute to the Rayleigh wave excitation with 0 and 90 degrees IDTs. Similar interpretations can be drawn by merely looking at the rotated piezoelectric tensor. As a reminder, the piezoelectric tensor for tetragonal c-axis (Ti atom is out of the wafer’s plane) SrTiO3 is given by
$e]=\begin{bmatrix}0&0&0&0&e_{15}&0\\ 0&0&0&e_{15}&0&0\\ e_{31}&e_{31}&e_{33}&0&0&0\end{bmatrix}$. (S10)
@@
$e]_{a}^{0}=\begin{array}[]{cccccc}S_{1}\quad S_{2}\quad S_{3}\quad S_{4}\quad S_{5}\quad S_{6}\\ \left[\begin{array}[]{cccccc}0&0&0&0&0&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{15}\\ e_{31}&e_{33}&e_{31}&0&0&0\\ 0&0&0&e_{15}&0&0\\ \end{array}\right]\\ \end{array}\begin{gathered}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}E_{x}\\ E_{y}\\ E_{z}\end{gathered}$ (S11)
@@
$e]_{c}^{90}=\begin{array}[]{cccccc}S_{1}\quad S_{2}\quad S_{3}\quad S_{4}\quad S_{5}\quad S_{6}\\ \left[\begin{array}[]{cccccc}\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{33}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{31}&\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}e_{31}&0&0&0\\ 0&0&0&0&0&e_{15}\\ 0&0&0&0&e_{15}&0\\ \end{array}\right]\\ \end{array}\begin{gathered}\\ \color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}E_{x}\\ E_{y}\\ E_{z}\end{gathered}$ (S12)
For a dominantly x-oriented electric field across $e]_{a}^{0}$ the $e_{15}$ now is at 1,6 position which couples x-field to shear x-y stress, consistent with the mode profile shown in Fig. S5a. Similarly, for x-oriented electric field across $e]_{c}^{90}$, the $e_{33}$ now is at 1,1 position which couples x-field to x -longitudinal stress, and $e_{31}$ at the 1,3 position which couples x-field to z- transverse stresses. This matches the elliptical motion that builds the Rayleigh wave.
Fig. S5: Simulated acoustic mode profiles. Simulated mode profiles with its corresponding driving domain for (a) Rayleigh wave and (b) Bulk shear wave.
@@ 1.5 Single domain piezoelectric coefficient of twin-domain structure
Starting from the electromechanical coupling coefficient definition (?)
$K=\frac{U_{m}}{\sqrt{U_{e}U_{d}}},$ (S13)
@@
$U_{m}=\frac{1}{2}\left[\int_{\Omega_{a}}\left(\frac{e_{33}^{(a)}}{c_{33}^{E(a)}}\right)E_{3}T_{3}dV+\int_{\Omega_{c}}\left(\frac{e_{33}^{(c)}}{c_{33}^{E(c)}}\right)E_{3}T_{3}dV\right],$ (S14)
$U_{e}$ $=\frac{1}{2}\int_{\Omega}\frac{1}{c_{33}^{E}}T_{3}^{2}dV,$ (S15)
$U_{d}$ $=\frac{1}{2}\int_{\Omega}\varepsilon_{33}^{T}E_{3}^{2}dV.$ (S16)
@@
$\operatorname{Vol}\left(\Omega_{a}\right)=\operatorname{Vol}\left(\Omega_{c}\right)=\frac{1}{2}\operatorname{Vol}(\Omega).$ (S17)
@@
$U_{m}=\frac{1}{2}\left[\int_{\Omega_{x}}\left(\frac{e_{15}^{(x)}}{c_{55}^{E(x)}}\right)\cos(45^{\circ})^{2}E_{1}T_{5}dV+\int_{\Omega_{y}}\left(\frac{e_{15}^{(y)}}{c_{55}^{E(y)}}\right)\cos(45^{\circ})^{2}E_{1}T_{5}dV\right].$ (S19)
@@
$U_{e}$ $=\frac{1}{2}\int_{\Omega}\frac{1}{c_{55}^{E}}T_{5}^{2}dV,$ (S20)
$U_{d}$ $=\frac{1}{2}\int_{\Omega}\varepsilon_{11}^{T}E_{1}^{2}dV.$ (S21)